# Panning for Gold: Expanding Domain-Specific Knowledge Graphs with General Knowledge

### v1

@@ 5. Experiments
    Can ExeFuse effectively address the challenges of practical challenges and achieve significant improvements? (Section 5.2)
@@
    Does the new task offer sufficient research value, and do the DKGF(W-I) and DKGF(Y-I) datasets provide meaningful insights for the solution to this new task? (Section 5.2)
@@
    How much do each module in the ExeFuse contribute to the final results? Are these modules reasonably designed? (Section 5.3)
@@
    Does the ExeFuse successfully balance alignment accuracy and efficiency? (Section 5.4)
@@
    Which aspects of this challenging task require further in-depth investigation? (Section 5.5)
@@
    Are there intuitive cases to straightly demonstrate the effectiveness of ExeFuse? (Section 5.6)
@@ 5.1.1. DKGF Benchmarks
Currently, there is no benchmark that effectively studies and evaluates the DKGF task. To this end, we construct two new DKGF benchmarks, i.e., DKGF(W-I) and DKGF(Y-I), designed to support knowledge integration and application across domains. As shown in Table 1, the domain-specific knowledge is drawn from the Integrated Crisis Early Warning System (ICEWS) database, which provides data on temporal political events that capture evolving geopolitical relationships among political actors. General knowledge comes from Wikidata and YAGO, both of which are high-quality, large-coverage resources that provide rich background facts and entity attributes. During dataset construction, we align and fuse cross-domain knowledge by incorporating relevant information from Wikidata and YAGO into ICEWS. This process mitigates the incompleteness of domain knowledge, improves multidimensional event understanding, and offers more comprehensive data support for knowledge reasoning and applications in integrated crisis early warning tasks. The detailed construction processes are illustrated in the technical report333 https://github.com/eduzrh/DKGF_benchmark/blob/main/VLDB_technical_report.pdf.
5.1.2. Benchmark Configurations
Currently, there is no specific solution for the DKGF task. To establish a comprehensive evaluation framework for DKGF, we systematically propose 21 representative and classic benchmark configurations for extensive comparison. These configurations can be divided into general-purpose configurations and cross-task adaptation configurations. More details can be found in the technical report3.
@@ 5.1.2. Benchmark Configurations
    General-purpose Configurations. To establish benchmark performance using universal architectures, this category employs broadly applicable methods (e.g., TransE, GNN, BERT, LLMs) rather than task-specific designs. This category of configurations is mainly based on current advanced and classic general methods, including rule-based methods (i.e., ”Rule.”), such as StringMatch-F (Du et al., 2025) and TF-IDF-F (Paulsen et al., 2023); Translation-based methods (i.e., “Trans.”) such as TransE-F (Bordes et al., 2013), TransH-F (Sun et al., 2019), DistMult-F (Yang et al., 2015), and ComplEx-F (Trouillon et al., 2016); GNN-based methods (i.e., “GNN”) like GCN-F (Chen et al., 2024c; Zhang et al., 2022), TransGNN-F (Zhang et al., 2024b), and Graph-Memba-F (Behrouz and Hashemi, 2024); Generative methods (i.e., “Generative.”) like BERT-F (Devlin et al., 2019a), ICL-F (Kothapalli et al., 2025; Wei et al., 2022), Self-Consistency-F (Wang et al., 2023b; Pan et al., 2024b), and Self-RAG-F (Asai et al., 2024);
@@
    Cross-task Adaptation Configurations. This category of configuration mainly involves improving representative methods from current related research tasks to adapt to the DKGF task, including entity alignment (i.e., “EA.”), such as SimpleHHEA-F (Jiang et al., 2024c) and ChatEA-F (Jiang et al., 2024b); knowledge graph completion (i.e., “KGC.”), such as KG-BERT-F (Yao et al., 2019b; Zhang et al., 2024d), KG-LLaMA-F (Yao et al., 2025; Zhang et al., 2024d), KoPA-F (Zhang et al., 2024d) and PRGC-F (Lv et al., 2022b); and relation triple extraction (i.e., “RTE.”), like NoGen-BART-F (Li et al., 2024b), and NoGen-T5-F (Li et al., 2024b).
@@ 5.1.3. Evaluation Protocol and Implementation Details
To evaluate the quality of DKGF produced by ExeFuse and the benchmark configurations, we follow the triple classification task (Li et al., 2024b; Zhang et al., 2024d) and use accuracy (ACC), precision (P), recall (R), and F1-score (F1) as evaluation metrics. This task is essentially a binary classification problem, and all test datasets are label-balanced. We also follow two classic data split settings (Zhang et al., 2024d; Jiang et al., 2024b), where DKGF(W-I)-S1 and DKGF(Y-I)-S1 use 80% of training data, and DKGF(W-I)-S2 and DKGF(Y-I)-S2 use 70%.
Additionally, we use runtime (in seconds) as an efficiency metric. All LLMs are implemented using the same model version, GPT-4 (gpt-4-0125-preview). We implement ExeFuse in PyTorch with embedding dimension $d=100$. The model uses 20 semantic operators organized into three granularity levels and is trained for 50 epochs using the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a batch size of 512. Negative samples are dynamically generated with a 1:1 ratio to positive facts. We optimize the fusion probability in Eq. (10) using binary cross-entropy or Focal Loss. During evaluation, a general fact is regarded as domain-relevant if its fusion probability is jointly determined by program executability in Eq. (8) and structural consistency in Eq. (9) and exceeds a fixed threshold of 0.5. Hyperparameters for all configurations are tuned in the validation set using a grid search, following the ranges recommended in the original papers. More details can be found in the technical report3.
Benchmark Configurations Settings DKGF(W-I)-S1 DKGF(Y-I)-S1 DKGF(W-I)-S2 DKGF(Y-I)-S2 Sema. Struc. LLM. ACC P R F1 ACC P R F1 ACC P R F1 ACC P R F1 General-purpose Rule. StringMatch-F 0.496 0.160 0.002 0.004 0.498 0.286 0.003 0.005 0.498 0.192 0.002 0.003 0.498 0.308 0.003 0.005 TF-IDF-F 0.498 0.461 0.025 0.047 0.507 0.598 0.043 0.081 0.498 0.451 0.023 0.043 0.508 0.598 0.046 0.086 Trans. TransE-F 0.649 0.668 0.596 0.630 0.588 0.782 0.244 0.372 0.637 0.662 0.560 0.607 0.575 0.772 0.212 0.332 TransH-F 0.640 0.683 0.520 0.591 0.583 0.764 0.240 0.365 0.632 0.685 0.489 0.570 0.568 0.768 0.195 0.311 DistMult-F 0.554 0.547 0.629 0.585 0.525 0.524 0.548 0.536 0.536 0.532 0.590 0.560 0.532 0.530 0.562 0.545 ComplEx-F 0.553 0.546 0.632 0.586 0.526 0.525 0.567 0.545 0.544 0.539 0.609 0.572 0.515 0.514 0.546 0.530 GNN. GCN-F 0.485 0.487 0.572 0.526 0.501 0.501 0.793 0.614 0.493 0.494 0.582 0.534 0.498 0.499 0.796 0.613 TransGNN-F 0.490 0.491 0.505 0.498 0.465 0.465 0.477 0.471 0.486 0.487 0.507 0.497 0.468 0.468 0.481 0.474 Graph-Memba-F 0.485 0.487 0.572 0.526 0.503 0.502 0.794 0.615 0.493 0.494 0.582 0.534 0.499 0.501 0.796 0.615 Generative. BERT-F 0.532 0.531 0.549 0.540 0.586 0.599 0.523 0.558 0.530 0.531 0.507 0.519 0.569 0.561 0.631 0.594 ICL-F 0.550 0.553 0.528 0.540 0.583 0.573 0.649 0.609 0.494 0.494 0.523 0.508 0.488 0.489 0.523 0.505 Self-Consistency-F 0.592 0.607 0.522 0.561 0.590 0.584 0.621 0.602 0.531 0.533 0.507 0.520 0.566 0.559 0.628 0.591 Self-RAG-F 0.576 0.589 0.506 0.544 0.594 0.623 0.477 0.540 0.544 0.544 0.545 0.545 0.577 0.595 0.481 0.532 Cross-task Adaptation EA. SimpleHHEA-F 0.490 0.492 0.507 0.499 0.493 0.493 0.503 0.498 0.492 0.492 0.507 0.500 0.477 0.477 0.481 0.479 ChatEA-F 0.596 0.551 0.581 0.566 0.649 0.727 0.477 0.576 0.592 0.611 0.507 0.554 0.592 0.619 0.481 0.541 KGC. KG-BERT-F 0.523 0.522 0.546 0.534 0.516 0.515 0.531 0.523 0.507 0.507 0.515 0.511 0.503 0.503 0.481 0.492 KG-LLaMA-F 0.512 0.512 0.531 0.521 0.539 0.537 0.558 0.547 0.529 0.527 0.529 0.528 0.521 0.521 0.514 0.517 KoPA-F 0.558 0.556 0.577 0.566 0.561 0.556 0.602 0.578 0.549 0.547 0.564 0.556 0.541 0.543 0.522 0.532 PRGC-F 0.503 0.503 0.508 0.505 0.507 0.507 0.515 0.511 0.504 0.504 0.507 0.506 0.503 0.503 0.481 0.492 RTE. NoGen-BART-F 0.516 0.515 0.531 0.523 0.511 0.511 0.523 0.517 0.508 0.508 0.517 0.512 0.509 0.509 0.491 0.500 NoGen-T5-F 0.510 0.510 0.522 0.516 0.507 0.507 0.515 0.511 0.509 0.509 0.519 0.514 0.504 0.504 0.509 0.506 ExeFuse (Ours) 0.680 0.631 0.669 0.690 0.661 0.615 0.717 0.662 0.655 0.627 0.683 0.654 0.633 0.634 0.682 0.657
@@ 5.2.1. Comparison with All Configurations
As shown in Table 2, to address RQ1 and RQ2, we have performed a comprehensive comparison for this new task. These results lead to several key observations:
@@
    Overall Observation. Across all configurations, ExeFuse achieves up to 9.5% relative improvement in ACC and F1 over the strongest baselines. These results demonstrate that DKGF requires explicitly modeling domain relevance and cross-domain abstraction gaps, which are not adequately addressed by existing general-purpose or cross-task adaptation configurations.
@@
    Evaluation of Task Value and Dataset Challenge. Furthermore, we perform a comprehensive analysis of the DKGF(W-I) and DKGF(Y-I) datasets, focusing on their complexity of fusion and practical utility. As shown in Table 2, the results indicate a significant potential for improvement in overall performance, as evidenced by the F1 scores of the state-of-the-art benchmark configurations, which are only 0.607 and 0.615 for DKGF(W-I)-S2 and DKGF(Y-I)-S2, respectively. It is worth noting that the performance of most models in related research tasks, such as entity alignment and knowledge graph completion, exceeds 0.9 in many cases (Jiang et al., 2024b; Zhang et al., 2024d). These findings highlight the inherent difficulty of the task and suggest that there are still numerous unresolved challenges.
@@ 5.2.2. Transferability Exploration
As shown in Figure 3, we demonstrate the transferability by evaluating the performance of ExeFuse on entities that do not appear during training. Specifically, we partition the test set facts into two categories: seen (S) and unseen (U). Specifically, the facts in the test set are divided into two parts: the seen (S) part and the unseen (U) part. If the head or tail entity in a fact is not present in the training set, it is regarded as unseen. In this setting, a subset of entities does not participate in the training process, making it difficult for the model to locate these truly relevant entities, which makes the test process more challenging. We report the accuracy and F1 scores for the test facts seen (S), unseen (U), and all (A) in four key configurations: ExeFuse, DistMult-F, Self-RAG-F, and ChatEA-F.
The experimental results reveal that ExeFuse outperforms the other methods in unseen entities. In contrast, the performance of current general-purpose and cross-task adaptation benchmarks exhibits considerable instability. This indicates that ExeFuse does not rely on entity memorization, but instead learns transferable relevance patterns grounded in domain-level execution behavior on the DKG. Notably, this learned mapping retains its transferability even when entities remain unseen during training.
5.3. Ablation Study (RQ3)
To address RQ3, we conduct an ablation analysis to investigate the effect of ExeFuse. As shown in Table 3, we can observe that: ① By removing the Granularity-Aware Operator Mapping, the performance drops by 11.8%, indicating that the three-level granularity hierarchy is essential for addressing Challenge II (Cross-Domain Knowledge Granularity Misalignment). ② Removing Stage 1 (operator mapping and program composition) causes the largest performance drop of 14.7%, confirming that granularity-aware program preparation is the most critical component. ③ Removing Stage 2 (Program Execution) results in a 13.2% drop, showing that executing programs on the DKG structure is necessary for adapting general knowledge to domain contexts. ④ Removing Stage 3 (Executability Verification) causes a 10.3% drop, demonstrating that the Relevance = Executability mechanism is crucial for addressing Challenge I (High Ambiguity of Domain Relevance). ⑤ By including all stages, ExeFuse achieves the best performance, indicating that all three stages work synergistically to address both challenges.
Figure 3. The results of the transferability experiment on the DKGF(W-I) and DKGF(Y-I) datasets at different training rates. The test data is split into seen (S) and unseen (U) parts based on whether the entity appeared during training. The total results for all (A) test data are also reported. Accuracy (ACC) and F1-score (F1) are shown in the radar charts.
5.4. Efficiency Analysis (RQ4)
To comprehensively address RQ4, we evaluated our framework in terms of both computational efficiency and token consumption, comparing it with two representative categories of configurations: ① lightweight general-purpose and cross-task adaptation configurations (e.g., TransH-F, SimpleHHEA-F), which are efficient but limited in capturing complex domain interactions, and ② LLM-based general-purpose and cross-task adaptation configurations (e.g., ChatEA-F, Self-Consistency-F), which exploit LLMs but incur high computational overhead. As shown in Table 4 , ExeFuse achieves the best trade-off between effectiveness and efficiency: it delivers the highest accuracy (Avg. 0.685) while eliminating token consumption through our small-model paradigm, requiring only 42.6 seconds with substantial performance gains (17.9%-38.4%). This validates the advantage of our Fact-as-Program paradigm, which provides a new perspective for DKGF by reformulating knowledge fusion as executable semantic reasoning without relying on expensive LLM API calls. These results validate its robustness in overcoming performance bottlenecks on the DKGF benchmark.
Table 3. The results of ablation study. “ACC”: accuracy. “F1”: F1 score. “Avg.”: average performance across the accuracy and F1 score metrics. $\Delta$”: the performance drop compared to the full model.
@@ 5.4. Efficiency Analysis (RQ4)
ExeFuse (Full) 0.680 0.690 0.685 -
ExeFuse w/o Domain Fact Generator 0.620 0.623 0.622 -9.2%
ExeFuse w/o Granularity-Aware Mapping 0.595 0.601 0.598 -12.7%
ExeFuse w/o Stage 1 0.570 0.582 0.576 -15.9%
ExeFuse w/o Stage 2 0.577 0.583 0.580 -15.3%
ExeFuse w/o Stage 3 0.610 0.622 0.616 -10.1%
Table 4. Accuracy and cost analysis of ExeFuse and benchmark configurations on DKGF(W-I)-S1. Avg.tokens: the average tokens cost of the model. Avg.time: the average time (seconds) cost of the model.

### v2

@@ 4.7. Theoretical Motivation
Benchmark Configurations Settings DKGF(W-I)-S1 DKGF(Y-I)-S1 DKGF(W-I)-S2 DKGF(Y-I)-S2 Sema. Struc. LLM. ACC P R F1 ACC P R F1 ACC P R F1 ACC P R F1 General-purpose Rule. StringMatch-F 0.496 0.160 0.002 0.004 0.498 0.286 0.003 0.005 0.498 0.192 0.002 0.003 0.498 0.308 0.003 0.005 TF-IDF-F 0.498 0.461 0.025 0.047 0.507 0.598 0.043 0.081 0.498 0.451 0.023 0.043 0.508 0.598 0.046 0.086 Trans. TransE-F 0.649 0.668 0.596 0.630 0.588 0.782 0.244 0.372 0.637 0.662 0.560 0.607 0.575 0.772 0.212 0.332 TransH-F 0.640 0.683 0.520 0.591 0.583 0.764 0.240 0.365 0.632 0.685 0.489 0.570 0.568 0.768 0.195 0.311 DistMult-F 0.554 0.547 0.629 0.585 0.525 0.524 0.548 0.536 0.536 0.532 0.590 0.560 0.532 0.530 0.562 0.545 ComplEx-F 0.553 0.546 0.632 0.586 0.526 0.525 0.567 0.545 0.544 0.539 0.609 0.572 0.515 0.514 0.546 0.530 GNN. GCN-F 0.485 0.487 0.572 0.526 0.501 0.501 0.793 0.614 0.493 0.494 0.582 0.534 0.498 0.499 0.796 0.613 TransGNN-F 0.490 0.491 0.505 0.498 0.465 0.465 0.477 0.471 0.486 0.487 0.507 0.497 0.468 0.468 0.481 0.474 Graph-Memba-F 0.485 0.487 0.572 0.526 0.503 0.502 0.794 0.615 0.493 0.494 0.582 0.534 0.499 0.501 0.796 0.615 Generative. BERT-F 0.532 0.531 0.549 0.540 0.586 0.599 0.523 0.558 0.530 0.531 0.507 0.519 0.569 0.561 0.631 0.594 ICL-F 0.550 0.553 0.528 0.540 0.583 0.573 0.649 0.609 0.494 0.494 0.523 0.508 0.488 0.489 0.523 0.505 Self-Consistency-F 0.592 0.607 0.522 0.561 0.590 0.584 0.621 0.602 0.531 0.533 0.507 0.520 0.566 0.559 0.628 0.591 Self-RAG-F 0.576 0.589 0.506 0.544 0.594 0.623 0.477 0.540 0.544 0.544 0.545 0.545 0.577 0.595 0.481 0.532 Cross-task Adaptation EA. SimpleHHEA-F 0.490 0.492 0.507 0.499 0.493 0.493 0.503 0.498 0.492 0.492 0.507 0.500 0.477 0.477 0.481 0.479 ChatEA-F 0.596 0.551 0.581 0.566 0.649 0.727 0.477 0.576 0.592 0.611 0.507 0.554 0.592 0.619 0.481 0.541 KGC. KG-BERT-F 0.523 0.522 0.546 0.534 0.516 0.515 0.531 0.523 0.507 0.507 0.515 0.511 0.503 0.503 0.481 0.492 KG-LLaMA-F 0.512 0.512 0.531 0.521 0.539 0.537 0.558 0.547 0.529 0.527 0.529 0.528 0.521 0.521 0.514 0.517 KoPA-F 0.558 0.556 0.577 0.566 0.561 0.556 0.602 0.578 0.549 0.547 0.564 0.556 0.541 0.543 0.522 0.532 PRGC-F 0.503 0.503 0.508 0.505 0.507 0.507 0.515 0.511 0.504 0.504 0.507 0.506 0.503 0.503 0.481 0.492 RTE. NoGen-BART-F 0.516 0.515 0.531 0.523 0.511 0.511 0.523 0.517 0.508 0.508 0.517 0.512 0.509 0.509 0.491 0.500 NoGen-T5-F 0.510 0.510 0.522 0.516 0.507 0.507 0.515 0.511 0.509 0.509 0.519 0.514 0.504 0.504 0.509 0.506 ExeFuse (Ours) 0.680 0.673 0.708 0.690 0.661 0.615 0.717 0.662 0.655 0.627 0.683 0.654 0.633 0.634 0.682 0.657
Table 3. Transferability analysis on DKGF(W-I)-S1.
Model Seen Unseen All
ACC F1 ACC F1 ACC F1
DistMult-F 0.563 0.623 0.491 0.340 0.554 0.585
Self-RAG-F 0.565 0.570 0.594 0.372 0.576 0.544
ChatEA-F 0.615 0.610 0.675 0.429 0.632 0.589
ExeFuse (Ours) 0.672 0.709 0.731 0.565 0.680 0.690
@@ 5. Experiments
    Can ExeFuse effectively address the challenges of challenges and achieve significant improvements?
@@
    Does the new task offer sufficient research value, and do the DKGF(W-I) and DKGF(Y-I) datasets provide meaningful insights for the solution to this new task?
@@
    How much do each module in the ExeFuse contribute to the final results? Are these modules reasonably designed?
@@
    Does the ExeFuse successfully balance alignment accuracy and efficiency?
@@
    Which aspects of this challenging task require further in-depth investigation?
@@
    Are there intuitive cases to straightly demonstrate the effectiveness of ExeFuse?
@@ 5.1.1. DKGF Benchmarks
Currently, there is no benchmark that effectively studies and evaluates the DKGF task. To this end, we construct two new DKGF benchmarks, i.e., DKGF(W-I) and DKGF(Y-I), designed to support knowledge integration and application across domains. As shown in Table 1, the DKG is drawn from the Integrated Crisis Early Warning System (ICEWS) database, which provides data on temporal political events that capture evolving geopolitical relationships among political actors. General knowledge comes from Wikidata and YAGO, both of which are high-quality, large-coverage resources that provide rich background facts and entity attributes. During dataset construction, we align and fuse cross-domain knowledge by incorporating relevant information from Wikidata and YAGO into ICEWS. This process mitigates the incompleteness of domain knowledge, improves multidimensional event understanding, and offers more comprehensive data support for knowledge reasoning and applications in integrated crisis early warning tasks.
5.1.2. Benchmark Suite Construction & Adaptation
Since DKGF is a nascent task with no pre-existing specialized solutions, a primary contribution of this work is to establish a comprehensive and standardized evaluation suite to facilitate future research. We systematically adapted 21 representative methods from related fields (e.g., KGC, Entity Alignment) to the DKGF setting using a unified Fusion Scoring Protocol. Specifically, to bridge the gap between their original tasks and DKGF: (1) for translation and GNN-based methods (e.g., TransE, GCN), we retrained them on the joint graph $G^{d}\cup G^{g}$ to score candidate fusion triples $(h^{d},r^{d},t^{g})$ based on link plausibility; (2) for Entity Alignment methods (e.g., ChatEA), we utilized aligned entity pairs to project potential facts from $G^{g}$ into the $G^{d}$ semantic space. This rigorous adaptation not only ensures a fair comparison but also provides a robust and reproducible benchmark configuration ecosystem for the community. These configurations are categorized as follows:
@@ 5.1.2. Benchmark Suite Construction & Adaptation
    General-purpose Configurations. To establish benchmark performance using universal architectures, this category employs broadly applicable methods (e.g., TransE, GNN, BERT, LLMs) rather than task-specific designs. This category of configurations is mainly based on current advanced and classic general methods, including rule-based methods (i.e., “Rule.”), such as StringMatch-F (Du et al., 2025) and TF-IDF-F (Paulsen et al., 2023); Translation-based methods (i.e., “Trans.”) such as TransE-F (Bordes et al., 2013), TransH-F (Sun et al., 2019), DistMult-F (Yang et al., 2015), and ComplEx-F (Trouillon et al., 2016); GNN-based methods (i.e., “GNN”) like GCN-F (Chen et al., 2024c; Zhang et al., 2022), TransGNN-F (Zhang et al., 2024b), and Graph-Memba-F (Behrouz and Hashemi, 2024); Generative methods (i.e., “Generative.”) like BERT-F (Devlin et al., 2019a), ICL-F (Kothapalli et al., 2025; Wei et al., 2022), Self-Consistency-F (Wang et al., 2023b; Pan et al., 2024b), and Self-RAG-F (Asai et al., 2024);
@@
    Cross-task Adaptation Configurations. This category of configuration mainly involves improving representative methods from current related research tasks to adapt to the DKGF task, including entity alignment (i.e., “EA.”), such as SimpleHHEA-F (Jiang et al., 2024b) and ChatEA-F (Jiang et al., 2024a); knowledge graph completion (i.e., “KGC.”), such as KG-BERT-F (Yao et al., 2019b; Zhang et al., 2024d), KG-LLaMA-F (Yao et al., 2025; Zhang et al., 2024d), KoPA-F (Zhang et al., 2024d) and PRGC-F (Lv et al., 2022b); and relation triple extraction (i.e., “RTE.”), like NoGen-BART-F (Li et al., 2024b), and NoGen-T5-F (Li et al., 2024b).
Baseline adaptation protocol.
Adapting existing methods to DKGF is non-trivial, as the task involves cross-graph fusion without shared embedding spaces. For each baseline, we follow its strongest feasible adaptation under a unified information setting, without introducing task-specific heuristics unavailable to other methods. For example, translation-based KGC models (e.g., TransE) are trained on the unified graph $G^{d}\cup G^{g}$ with disjoint entity embeddings, and candidate triples $(h^{d},r^{d},t^{g})$ are scored using their standard plausibility functions. Entity alignment methods leverage aligned entity pairs to project candidate facts, while RTE-based methods operate purely on structured triples without access to gold fusion labels.
@@ 5.1.3. Evaluation Protocol and Implementation Details
To evaluate the quality of DKGF produced by ExeFuse and the benchmark configurations, we follow the triple classification task (Li et al., 2024b; Zhang et al., 2024d) and use accuracy (ACC), precision (P), recall (R), and F1-score (F1) as evaluation metrics. This task is essentially a binary classification problem, and all test datasets are label-balanced. We also follow two classic data split settings (Zhang et al., 2024d; Jiang et al., 2024a), where DKGF(W-I)-S1 and DKGF(Y-I)-S1 use 80% of training data, and DKGF(W-I)-S2 and DKGF(Y-I)-S2 use 70% to test robustness. We use runtime (in seconds) as an efficiency metric. All LLMs are implemented using the same model version, GPT-4 (gpt-4-0125-preview).
@@ 5.2.1. Comparison with All Configurations
As shown in Table 2, we have performed a comprehensive comparison for this new task. These results lead to several key observations:
@@
    Overall Observation. Across all configurations, ExeFuse achieves up to 9.5% relative improvement in ACC and F1 over the strongest benchmark configurations. These results demonstrate that DKGF requires explicitly modeling domain relevance and cross-domain abstraction gaps, which are not adequately addressed by existing general-purpose or cross-task adaptation configurations.
@@
    Evaluation of Task Value and Dataset Challenge. Furthermore, we perform a comprehensive analysis of the DKGF(W-I) and DKGF(Y-I) datasets, focusing on their complexity of fusion and practical utility. As shown in Table 2, the results indicate a significant potential for improvement in overall performance, as evidenced by the F1 scores of the state-of-the-art benchmark configurations, which are only 0.607 and 0.615 for DKGF(W-I)-S2 and DKGF(Y-I)-S2, respectively. It is worth noting that the performance of most models in related research tasks, such as entity alignment and knowledge graph completion, exceeds 0.9 in many cases (Jiang et al., 2024a; Zhang et al., 2024d). These findings highlight the inherent difficulty of the task and suggest that there are still numerous unresolved challenges.
@@ 5.2.2. Transferability Exploration
To verify that ExeFuse learns transferable logic rather than memorizing entities, we evaluated performance on unseen entities. Specifically, we partition the test set facts into two categories: seen and unseen. Specifically, the facts in the test set are divided into two parts: the seen part and the unseen part. If the head or tail entity in a fact is not present in the training set, it is regarded as unseen. In this setting, a subset of entities does not participate in the training process, making it difficult for the model to locate these truly relevant entities, which makes the test process more challenging. As shown in Table 3, ExeFuse significantly outperforms baselines on unseen data, maintaining consistently high F1 scores. In contrast, methods like ChatEA-F exhibit instability when facing entities not present during training. This confirms that our neuro-symbolic execution paradigm generalizes to new geometric regions of the DKG.
Table 4. Ablation study and design verification on DKGF(W-I)-S1. “Avg.”: average of ACC and F1. $\Delta$”: relative performance drop in Avg.
@@
ExeFuse (Full Model) 0.680 0.690 0.685 -
I. Complexity Validation (Is the complex pipeline necessary?)
w/o Rules $\mathcal{L}$ (Single Affine Projection) 0.605 0.612 0.609 -11.1%
w/o Reasoning (Direct Classification) 0.577 0.583 0.580 -15.3%
II. Module Necessity (Do all components contribute?)
w/o Structure-Aware Encoder ($\phi_{enc}$) 0.570 0.582 0.576 -15.9%
w/o Neural Rule Selector ($p(R|q^{g})$) 0.620 0.623 0.622 -9.2%
w/o Granularity Calibration ($S_{exec}$) 0.595 0.601 0.598 -12.7%
III. Design Choice Verification (Are choices optimal?)
Replace Affine Rules with MLP 0.665 0.672 0.669 -2.3%
Replace KDE with Min-Distance 0.648 0.655 0.652 -4.8%
Table 5. Accuracy and cost analysis of ExeFuse and benchmark configurations on DKGF(W-I)-S1. Avg.tokens/Avg.time: the average tokens/time cost of the model.